OpenStudy (anonymous):

what is the square root of negative two over 9

OpenStudy (anonymous):

\[\sqrt{\frac{-2}{9}}\] \[\sqrt{\frac{-1*2}{3*3}}\] \[\sqrt{-1}=i\] \[\frac{i}{3}\sqrt{2}\]

OpenStudy (anonymous):

three square roots of 2 over 9

OpenStudy (anonymous):

or just square roots of two over 3

OpenStudy (anonymous):

well the negative is in the outside like this i will draw it:

OpenStudy (anonymous):

iam sorry i reas it wrong it looks like this i will draw it.

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

lololol!!